OpenStudy (anonymous):

(x-5)^2 = 49. Solve via method of square roots.

OpenStudy (cwrw238):

take square root of both sides: x - 5 = +/- 7

OpenStudy (zzr0ck3r):

take the square root of both sides.

OpenStudy (anonymous):

the answer in the packet says 12 or -2....how do i get there?

OpenStudy (cwrw238):

solve the 2 equations: x - 5 = 7 and x - 5 = -7

OpenStudy (anonymous):

OHHH thank you

OpenStudy (zzr0ck3r):

when you introduce a square root, you will have two answers (x-5)^2 = 49 take square root of both sides sqrt((x-5)^2) = sqrt(49) note: sqrt(49) has two solutions. namely 7,-7 (7*7 = 49, -7*-7 = 49) so we have x-5 = 7 AND x - 5 = -7 solving for BOTH we get x-5 = 7 implies x = 12 AND x-5 = -7 implies x = -2